An experiment is done on a horizontal surface. A b
Answer #1

a)

using conservation of energy

kinetic energy = spring potential energy

(0.5) m v2 = (0.5) k xo2

m1 v2 = k xo2

v = xo sqrt(k/m1)

b)

since the blocks sticks together after the collision , hence the collision is inelastic collision

using conservation of momentum

m1 v1i + m2 v2i = (m1 + m2) V

m1 xo sqrt(k/m1) + m2 (0) = (m1 + m2) V

V = xo sqrt(km1) /(m1 + m2)

c)

d = stopping distance

a = acceleration = - \muk g

Vf = final velocity = 0 m/s

Vi = initial velocity = xo sqrt(km1) /(m1 + m2)

using the equation

Vf2 = Vi2 + 2 a d

02 = x2o (km1)2 /(m1 + m2)2 + 2 (- \muk g) d

\muk = x2o (km1)2 /(m1 + m2)2 /(2gd)